How do I find the slope and y intercept?
X+3Y=-9

Answer by jai_kos(139)

Given x+ 3y = -9
The above equation can be rewritten as,
3y = - x - 9
y = -x/ 3 - 9 / 3
y = (-1/3) x - 3 ---->(1)
This is of the fiorm y = mx + c --->(2)
When you compare equation(1) with equation (2) we find that,
We find that, m = (-1/3) and c = (-3)
Here m represents the slope of the given equation and c represent the y intercept of the given equation.
Hence Slope = (-1/3) and y intercept = (-3)